# Assessing the quality of information extraction

### v1

@@ Conclusions
In this paper, we focused on quality evaluation of information extraction (IE) performed by large. First, we delved into the technical limitations of large language models (LLMs) complicating the extraction of information from a long context. To extract reasonable information from data it is needed to take into the account features such as context window limits, iterated extractions, extraction history recording and Lost in the middle phenomenon. Once the extraction is performed, assessing its quality is essential. However in many customized tasks, a truly objective method is missing, because of the lack of labeled data fitting the scope of the application. The versatile method presented in this paper overcomes the issue by adjustment of the data by insertion of an artificial information, a needle, into it. The artificial information created to this purpose is application and data-specific, but the method itself is applicable generally across the field of IE. By controlling the generation process of the needles, we created a synthetic ground truth that enables us to absolutely measure the extraction quality even when no manually labeled data is available. We introduced a MINEA score to measure the quality of extraction. The key part is a decision rule on whether a needle was successfully extracted or not. MINEA possibly combines several decision rules into one final score. Our empirical analysis of the MINEA score on a specialized dataset demonstrated its utility for evaluation of LLM-based IE tasks when ground truth is unavailable.
@@
Discovering a schema
Figure 4 shows a prompt to obtain the schema.org types from the attached text – Wikipedia article about IE [^5] https://en.wikipedia.org/wiki/Information_extraction. An LLM is asked to assign relevance to the types to distinguish the most important ones.

### v2

@@ Conclusions
In this paper, we focused on quality evaluation of information extraction (IE) performed by large language models (LLMs). First, we delved into the technical limitations of LLMs complicating the extraction of information from a long context. To extract reasonable information from data it is needed to take into the account features such as context window limits, iterated extractions, extraction history recording and Lost in the middle phenomenon. Once the extraction is performed, assessing its quality is essential. However in many customized tasks, a truly objective method is missing, because of the lack of labeled data fitting the scope of the application. The versatile method presented in this paper overcomes the issue by adjustment of the data by insertion of an artificial information, a needle, into it. The artificial information created to this purpose is application and data-specific, but the method itself is applicable generally across the field of IE. By controlling the generation process of the needles, we created a synthetic ground truth that enables us to absolutely measure the extraction quality even when no labeled data is available. We introduced a MINEA score to measure the quality of extraction. The key part is a decision rule on whether a needle was successfully extracted or not. MINEA possibly combines several decision rules into one final score. Our empirical analysis of the MINEA score on a specialized dataset demonstrated its utility for evaluation of LLM-based IE tasks when ground truth is unavailable.
@@
B1 Discovering a schema
Figure 4 shows a prompt to obtain the Schema.org types from the attached text – Wikipedia article about IE. [^5] https://en.wikipedia.org/wiki/Information_extraction An LLM is asked to assign relevance to the types to distinguish the most important ones.